An important part of the customer service responsibilities of a telephone company relate to the speed with which interruptions in residential service can be repaired on the same day. Suppose that past data indicate that the likelihood is .54 that interruptions in residential service can be repaired on the same day. For the first six interruptions reported on a given day, what is the probability that:
Answer #1

Solution : Given p = 0.54, n = 6

= > P(X = 6) = 6C6*0.54^6*0.46^0 = 0.0248

Binomial distribution = P(X = r) = nCr*p^r*q^(n-r)